OpenStudy (anonymous):

3(x + 1) = −2(x − 1) + 6

Parth (parthkohli):

Use distributive property to start with, and then isolate \(x\).

OpenStudy (anonymous):

3x+3=-2x+2+6 3x+2x=2+6-3 5x=5 x=5/5=1 hope u got it

OpenStudy (anonymous):

stebo plz give medal yarr

OpenStudy (theviper):

LOL i gave ;)

OpenStudy (anonymous):

ya thanxx..

Parth (parthkohli):

Nice, medals to direct answers.

Parth (parthkohli):

Thank you everybody! You are following our Code Of Conduct rather well! Giving medals to people who don't help! =) http://openstudy.com/code-of-conduct